40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  TCS elog  Not logged in ELOG logo
Message ID: 41     Entry time: Thu May 20 17:08:36 2010
Author: Aidan 
Type: Electronics 
Category: SLED 
Subject: SLED module - and driver - LIGO D1000892 - and Hartmann sensor 

Verified that the test-point for the current limit pot on the driver (Wavelength Electronics - LDTC 0520) was at 0.5V. Driver is set to INTERNAL set point at the moment. This is down about 10% below the current limited point.

Voltage across TP7 and TP9 = 0.970V = LD Current Mon

Voltage across TP2 and TP3 = 0.017V = LD P Mon

 

--- Hartmann sensor ---

-set the sampling rate on the CCD to 16HZ. With the current alignment and intensity this gives as maximum intensity of around 3850 out of 4095. Thus the pixels are not saturated.

- centroid_image located some of the spots - see attached image of spots where those located by the algorithm and circled. I need to play with the threshold level and spot_radius to get this to work properly.

 

 

Attachment 1: 2010-05-20_hartmann_image_and_located_spot.jpg  116 kB  Uploaded Thu May 20 18:39:39 2010  | Hide | Hide all
2010-05-20_hartmann_image_and_located_spot.jpg
ELOG V3.1.3-